OpenStudy (anonymous):

A coin is flipped 150 times. The results of the experiment are shown in the following table: Heads Tails 84 66 Which of the following statements best describes the experimental probability of getting heads? It is equal to the theoretical probability. It is 6% lower than the theoretical probability. It is 6% higher than the theoretical probability. The experimental probability cannot be concluded from the data in the table.

OpenStudy (anonymous):

@Jadeishere

OpenStudy (anonymous):

@uri

Vocaloid (vocaloid):

theoretical probability of getting heads is 0.5, or 50% experimental probability of getting heads is (number of heads)/(150) = 84/150 = 56%

OpenStudy (anonymous):

uhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

thank you lololol